# Reasons to Reject? Aligning Language Models with Judgments

### v1

@@ Abstract
As humans, we consistently engage in interactions with our peers and receive feedback in the form of natural language. This language feedback allows us to reflect on our actions, maintain appropriate behavior, and rectify our errors. The question arises naturally: can we use language feedback to align large language models (LLMs)? In contrast to previous research that aligns LLMs with reward or preference data, we present the first systematic exploration of alignment through the lens of language feedback (i.e., judgment). We commence with an in-depth investigation of potential methods that can be adapted for aligning LLMs with judgments, revealing that these methods are unable to fully capitalize on the judgments. To facilitate more effective utilization of judgments, we propose a novel framework, Contrastive Unlikelihood Training (CUT), that allows for fine-grained inappropriate content detection and correction based on judgments. Our offline alignment results show that, with merely 1317 off-the-shelf judgment data, CUT (LLaMA2-13b) can beat the 175B DaVinci003 and surpass the best baseline by 52.34 points on AlpacaEval. The online alignment results demonstrate that CUT can align LLMs (LLaMA2-chat-13b) in an iterative fashion using model-specific judgment data, with a steady performance improvement from 81.09 to 91.36 points on AlpacaEval. Our analysis further suggests that judgments exhibit greater potential than rewards for LLM alignment and warrant future research. [^1] Our source codes and models are publicly available at https://github.com/wwxu21/CUT
@@ 1 Introduction
Large language models (LLMs) acquire extensive knowledge and remarkable reasoning capabilities through self-supervised pre-training on large-scale corpora (Brown et al., 2020; Du et al., 2022; Touvron et al., 2023). To unleash the power of pre-trained LLMs for addressing real-world applications, it is crucial to ensure LLMs can follow human intentions and values (Ouyang et al., 2022). This process, known as alignment, has the potential to pave the way for a future in which artificial intelligence (AI) serves as a helpful and reliable partner for humanity (Wang et al., 2023b).
Figure 1 shows three typical paradigms to achieve alignment. The most straightforward one is learning from demonstrations, wherein human demonstrations of desired responses to a set of instructions are collected to fine-tune LLMs using supervised learning (Wei et al., 2022; Ouyang et al., 2022). However, the performance gains diminish rapidly when scaling up the data size (Zhou et al., 2023). On the other hand, learning from feedback offers a more scalable approach (Ouyang et al., 2022; Glaese et al., 2022; Bai et al., 2022a). One key advantage of feedback over demonstrations is that feedback can convey both positive and negative aspects, enabling the model to discern desirable and undesirable outcomes. In addition, feedback is usually tailored to the current model-generated responses, adhering to the principle of teaching according to the learner’s aptitude.
Prior research on learning from feedback has been primarily focused on incorporating value feedback (i.e., scalar rewards) into the alignment process. The most widely adopted reinforcement learning (RL) techniques, particularly proximal policy optimization (PPO) (Schulman et al., 2017; Ouyang et al., 2022), optimize an LLM to maximize the scalar rewards of its generated responses. Nevertheless, PPO is known to be complex and often unstable (Zheng et al., 2023), which has prompted numerous efforts to simplify or stabilize the training process (Ramamurthy et al., 2023; Peng et al., 2023b; Dong et al., 2023; Touvron et al., 2023). Another strand of work, referred to as Hindsight (Zhang et al., 2023; Liu et al., 2023a), transforms scalar rewards to language instructions and employs supervised training on the updated data.
@@
Language feedback (i.e., judgment) is another kind of feedback that provides more nuanced commendations and critiques through natural language expressions. Unlike scalar rewards, which are information-sparse for solely indicating the goodness of a response, judgments can elucidate the specific aspects that are good or bad, the rationale behind their evaluation, and suggestions for improvement. The above advantages suggest that aligning LLMs with judgments can be more efficient and effective (Saunders et al., 2022). However, current approaches merely use judgments to prompt LLMs for an improved response, which is subsequently employed as a new demonstration for supervised training (Bai et al., 2022b; Scheurer et al., 2022; 2023). This indirect utilization of judgments suffers from the incapability to learn from mistakes, which is the core spirit of learning from feedback, and is constrained by the refinement capabilities of LLMs.
@@
In this study, we present an extensive investigation of potential methods that can be adapted for aligning LLMs with judgments. To facilitate a comprehensive aligning process, we propose a novel framework, Contrastive Unlikelihood Training (CUT), that enables fine-grained inappropriate content detection and correction based on judgments. The core idea of CUT is to detect and penalize inappropriate content in a response by contrasting its generation probabilities guided by an authentic judgment that may contain negative opinions and a fabricated judgment portraying perfection.
@@
We carry out alignment experiments in both offline and online settings, wherein the target LLM learns from the off-the-shelf judgments and the judgments derived from self-generated responses, respectively. Extensive results on offline alignment demonstrate the effectiveness of CUT in learning from judgments in both cold-start (using unaligned base models such as LLaMA2) and warm-start (using aligned base models such as LLaMA2-chat) scenarios. Notably, when trained with only 1317 offline judgment data, CUT (LLaMA2-13b) attains a winning rate of 62.56 (beats the 175B DaVinci003222https://platform.openai.com/docs/model-index-for-researchers) and outperforms the best baseline by 52.34 points on AlpacaEval. Furthermore, our online alignment experiments show that CUT is capable of iteratively refining LLMs with up-to-date, model-specific judgments. For example, we observe a consistent performance improvement on LLaMA2-chat-13b over four times of CUT iterations, rising from 81.09 to 91.36 points on AlpacaEval. Our analysis comparing rewards and judgments suggests that aligning LLMs with judgments offers significant potential and warrants future research. Our contributions can be summarized as follows.
  • We present the first systematic exploration of aligning LLMs with judgments.
  • We introduce a novel framework, CUT, that facilitates the alignment of LLMs through direct and explicit learning from judgments. Notably, CUT allows fine-grained inappropriate content detection and correction based on judgments.
  • Our results showcase the effectiveness of CUT in aligning LLMs across cold-start and warm-start scenarios, generalist and specialist applications, as well as offline and online settings.
  • Our analysis indicates that judgments hold promising potential over rewards for aligning LLMs.
@@ 2 Related Work
2.1 Collecting Feedback
Value Feedback (Reward).
Traditional RL research for natural language processing (NLP) uses algorithmically defined metrics as reward functions, such as BLEU for translation (Wu et al., 2016) and ROUGE for summarization (Ranzato et al., 2016). For LLM alignment, existing works primarily leverage human preference data to fit a reward model, which subsequently generates scalar rewards (Ouyang et al., 2022). To augment the informativeness of value feedback, recent studies introduce rewards for multiple dimensions (Bai et al., 2022a; Touvron et al., 2023; Wu et al., 2023) and provide rewards for each sub-step (Lightman et al., 2023).
Language Feedback (Judgment).
Judgments typically necessitate human annotations on the model-generated responses. There are several works where judgments are collected for specific tasks, such as dialogue (Xu et al., 2023b), summarization (Saunders et al., 2022; Scheurer et al., 2022; 2023; Liu et al., 2023c), question answering (Li et al., 2022; Xu et al., 2023a), script generation (Tandon et al., 2022), and general instruction-following tasks (Wang et al., 2023a). Another direction is to train an AI judge to automatically provide precise judgments for the model’s responses (Bai et al., 2022b; Akyurek et al., 2023; Li et al., 2023).
2.2 Learning from Feedback
@@
Prompting.
Prompting does not alter the parameters of the LLMs. Instead, it leverages language feedback on previous responses to prompt the generation of improved responses (Welleck et al., 2022; Akyurek et al., 2023). The language feedback can be sourced from diverse aspects Nathani et al. (2023); Yu et al. (2023) and the refinement process can be iterated multiple times Yang et al. (2022); Peng et al. (2023a); Madaan et al. (2023). However, these approaches consume more computation than single-pass generation and usually rely on the in-context learning capabilities of the LLMs (Brown et al., 2020; Liu et al., 2023b).
Fine-tuning.
Fine-tuning aims to directly train a better LLM. In this context, value feedback has been extensively used through RL, particularly PPO (Schulman et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022a; Yang et al., 2023). However, these RL approaches are notoriously unstable and complex (Zheng et al., 2023). To stabilize RL, Ramamurthy et al. (2023) propose to reduce the action space through truncation and Peng et al. (2023b) employ an advantage model and selective rehearsal. In addition, many efforts have been put into designing simpler alternatives to RL. Dong et al. (2023); Touvron et al. (2023) treat value feedback as a ranking criterion and simply train models using the best model-generated responses. There are also attempts to leverage the results of prompting for training a better model. That is, the improved response elicited by language feedback is employed as new training data (Bai et al., 2022b; Scheurer et al., 2022; 2023). However, these methods still suffer from the incapability to learn from mistakes. Rafailov et al. (2023); Yuan et al. (2023); Song et al. (2023) demonstrate that LLMs themselves can be used as reward functions and derive different training objectives to eliminate the need for RL. Zhang et al. (2023); Liu et al. (2023a) relabel the input using the value feedback received by the response, referred to as Hindsight. This hindsight method allows LLMs to learn to generate responses of different qualities. In this work, our CUT is a novel fine-tuning method that allows LLMs to comprehensively learn from both positive and negative aspects of language feedback.
@@ 3 Preliminaries
In this section, we first lay out a formal problem definition of aligning LLMs with judgments and then present a survey of three potentially useful methods that can be adapted for tackling this problem.
@@ 3.1 Problem Setting
Suppose that there is a set of instruction-response-judgment triplets $(\bm{x},\bm{y},\bm{j})$, where the instruction $\bm{x}=[x_{1},\ldots,x_{M}]$, the response $\bm{y}=[y_{1},\ldots,y_{N}]$, and the judgment $\bm{j}=[j_{1},\ldots,j_{Q}]$ are token sequences of length $M$, $N$, and $Q$, respectively. The response may exhibit certain flaws or be considered entirely satisfactory. The judgment provides an analysis of the strengths and weaknesses of the response. The judgment can be drafted either by human annotators [^3] We argue that annotating judgments is not more difficult than annotating scalar rewards (or preferences), as annotators typically assign scalar rewards based on specific reasons. Essentially, we are just asking annotators to write down the reasons behind their decisions. or AI judge models (Akyurek et al., 2023; Li et al., 2023). The goal of aligning LLMs with judgments is to enable the LLM to retain appropriate behaviors mentioned in the strengths, and more importantly, address the weaknesses to prevent future misbehavior.
Depending on whether the responses $\bm{y}$ are generated from the LLM to be aligned, the learning process can be classified into two distinct types: offline alignment and online alignment. In offline alignment, the target LLM learns from a static, off-the-shelf, model-agnostic dataset. Conversely, in online alignment, the target LLM reflects on its own outputs through direct interactions with a judge. This online alignment process can be conducted iteratively, akin to how humans continuously improve their skills by receiving ongoing feedback from others over time.
3.2 Potential Solutions
Forward Prediction.
Forward prediction refers to the process of sequentially predicting the response and its judgment, which was originally proposed in the context of dialogue generation (Weston, 2016; Li et al., 2017). It can be seamlessly adapted to the alignment of LLMs. Specifically, the LLM is trained under the maximum likelihood estimation (MLE) objective to first generate the response $\bm{y}$ based on the instruction $\bm{x}$ and subsequently generate the judgment $\bm{j}$ based on the combined sequence $[\bm{x},\bm{y}]$.
$L^{f}(\bm{x},\bm{j},\bm{y})=-\frac{1}{N}\sum_{t}\log p_{\theta}(y_{t}|y_{<t},\bm{x})-\frac{1}{Q}\sum_{t}\log p_{\theta}(j_{t}|j_{<t},\bm{y},\bm{x})$ (1)
where $\theta$ represents the trainable parameters of the target LLM.
Imitation Learning from Language Feedback.
Imitation learning from Language Feedback (ILF) asks the LLM to refine the initial response $\bm{y}$ given the feedback $\bm{j}$ (Bai et al., 2022b; Scheurer et al., 2022; 2023). The improved response $\hat{\bm{y}}$, paired with the initial instruction $\bm{x}$, is used to fine-tune the LLM under the MLE objective.
$\hat{\bm{y}}$ $=\textbf{LLM}(\bm{x},\bm{y},\bm{j})$ (2)
$L^{i}(\bm{x},\hat{\bm{y}})$ $=-\frac{1}{N}\sum_{t}\log p_{\theta}(\hat{y}_{t}|\hat{y}_{<t},\bm{x})$ (3)
Hindsight.
Hindsight methods (Zhang et al., 2023; Liu et al., 2023a) rewrite the instruction $\bm{x}$ based on the scalar rewards received by the response $\bm{y}$. For instance, if a response receives a scalar reward below a certain threshold, the phrase “generate a correct answer” is appended to the original instruction; otherwise, “generate an incorrect answer” is added. Obviously, this approach can be naturally extended to our problem setting. Concretely, the LLM is trained to generate the response $\bm{y}$ conditioned on the sequence $[\bm{x},\bm{j}]$.
$L^{h}(\bm{x},\bm{j},\bm{y})=-\frac{1}{N}\sum_{t}\log p_{\theta}(y_{t}|y_{<t},\bm{x},\bm{j})$ (4)
However, in forward prediction, learning to generate judgments does not necessarily translate into enhanced response generation, given that response generation precedes judgment generation. ILF only makes use of the positive data (i.e., the improved responses), limiting the model’s capacity to recognize and rectify weaknesses or errors underscored in negative judgments. As for Hindsight, employing unsatisfactory responses as MLE targets inevitably increases the risk of generating unsatisfactory responses. In summary, we contend that existing methods cannot take full advantage of judgments, which motivates us to design a better solution.
4 Contrastive Unlikelihood Training
Instruction: $\bm{x}$ Response: $\bm{y}$ Judgment: $\bm{j}$ Aligned
$\bm{x}\xrightarrow{}\bm{y}$ $[\bm{x},\bm{j}]\xrightarrow{}\bm{y}$
@@ 4 Contrastive Unlikelihood Training
James had 4 packs of beef that were 5 pounds each. Each pack was 5 pounds and it cost 5.50. So 5 * 5.50 = 27.50 dollars.
@@
James had 4 packs of beef that were 5 pounds each. Each pack was 5 pounds and it cost 5.50. So 5 * 5.50 = 27.50 dollars.
@@
Table 1: The illustration of three categories of alignment data. The “Aligned” columns indicate if the response aligns with the instruction or the combination of instruction and judgment, respectively.
To overcome the limitations mentioned in § 3, we propose Contrastive Unlikelihood Training (CUT), a novel fine-tuning framework to align LLMs with judgments. The central idea of CUT can be summarized as Learning from Contrasting. We contrast the response generation under different conditions to shed light on the appropriate behavior that the LLM should maintain, as well as the specific content necessitating adjustments. Based on these insights, we use MLE training for appropriate content and Unlikelihood Training (UT) for inappropriate content.

### v2

@@ Abstract
As humans, we consistently interact with our peers and receive feedback in the form of natural language. This language feedback allows us to maintain appropriate behavior, and rectify potential errors. The question arises naturally: can we use language feedback to align large language models (LLMs)? In contrast to previous research that aligns LLMs with scalar rewards, we present the first systematic exploration of alignment through the lens of language feedback (i.e., judgment). We start with an in-depth investigation of potential methods that can be adapted for aligning LLMs with judgments, revealing that these methods cannot fully capitalize on judgments. To facilitate more effective utilization of judgments, we propose a novel framework, Contrastive Unlikelihood Training (CUT), that allows for fine-grained inappropriate content detection and correction based on judgments. Our results show that, with merely 1317 off-the-shelf judgment data, CUT (LLaMA2-13b) can beat the 175B DaVinci003 and surpass the best baseline by 48.51 points on AlpacaEval. CUT (LLaMA2-chat-13b) can also align LLMs in an iterative fashion using up-to-date model-specific judgments, improving performance from 81.09 to 91.68 points on AlpacaEval. Further analysis suggests that judgments hold greater potential than rewards in LLM alignment.
@@ 1 Introduction
Large language models (LLMs) acquire extensive knowledge and remarkable reasoning capabilities through large-scale pre-training (Brown et al., 2020; Du et al., 2022; Touvron et al., 2023). To unleash the power of pre-trained LLMs for addressing real-world applications, it is crucial to ensure LLMs can follow human values (Ouyang et al., 2022). This process, known as alignment, has the potential to pave the way for a future in which artificial intelligence (AI) serves as a helpful and reliable ally for humanity (Wang et al., 2023b).
Figure 1 shows three typical paradigms to achieve alignment. The most straightforward one is learning from demonstrations, wherein demonstrations of desired responses to a set of instructions are collected to fine-tune LLMs (Wei et al., 2022; Ouyang et al., 2022). However, the performance gains diminish rapidly when scaling up the data size (Zhou et al., 2023). On the other hand, learning from feedback offers a more scalable approach (Ouyang et al., 2022; Bai et al., 2022a). One key advantage of feedback over demonstrations is that feedback can convey both positive and negative aspects, enabling the model to discern desirable and undesirable outcomes. In addition, feedback is tailored to the current model, adhering to the principle of teaching according to the learner’s aptitude.
Prior research on learning from feedback primarily focuses on value feedback (i.e., scalar rewards). Reinforcement learning (RL) techniques, particularly PPO algorithm (Schulman et al., 2017), are employed to optimize an LLM to maximize the rewards of its generated responses. However, PPO is known to be complex and often unstable (Zheng et al., 2023), which has prompted numerous efforts to simplify or stabilize the training process (Ramamurthy et al., 2023; Peng et al., 2023b; Dong et al., 2023). Another strand of work, named Hindsight (Zhang et al., 2023; Liu et al., 2023a), transforms scalar rewards to language instructions and employs supervised training on the updated data.
@@
Language feedback (i.e., judgment) is another kind of feedback that offers nuanced commendations and critiques through natural language expressions. Unlike scalar rewards, which are information-sparse for solely indicating the goodness of a response, judgments can elucidate the specific aspects that are good or bad, the rationale behind their evaluation, and suggestions for improvement. The above advantages suggest that aligning LLMs with judgments can be more beneficial (Saunders et al., 2022). Current approaches merely use judgments to prompt LLMs for an improved response (Bai et al., 2022b; Scheurer et al., 2022; Yu et al., 2023). But this leaves the judgment minimally involved in the subsequent alignment stage. This indirect utilization of judgments suffers from the incapability to learn from mistakes, which is the core spirit of learning from feedback.
@@
In this study, we present an extensive investigation of potential methods that can be adapted for aligning LLMs with judgments. To facilitate a comprehensive aligning process, we propose a novel framework, Contrastive Unlikelihood Training (CUT), that enables fine-grained inappropriate content detection and correction based on judgments. CUT detects inappropriate content in a response by contrasting its generation probabilities under aligned and misaligned conditions and further penalizes the inappropriate content with unlikelihood training (UT) Welleck et al. (2020).
@@
We carry out experiments for both offline and online alignment, wherein the target LLM learns from the off-the-shelf judgments and the judgments derived from self-generated responses, respectively. Extensive results on offline alignment demonstrate the effectiveness of CUT in learning from judgments in both cold-start (using unaligned base LLMs such as LLaMA2) and warm-start (using aligned base LLMs such as LLaMA2-chat) scenarios. Notably, when trained with only 1317 offline judgment data, CUT (LLaMA2-13b) attains a winning rate of 61.06 and outperforms the best baseline by 48.51 points on AlpacaEval. Furthermore, our online alignment experiments show that CUT is capable of iteratively refining LLMs (LLaMA2-chat-13b) using model-specific judgments, with a steady performance improvement from 81.09 to 91.68 points on AlpacaEval. Our analysis comparing rewards and judgments suggests that aligning LLMs with judgments offers significant potential and warrants future research.
Our contributions can be summarized as follows: 1) We present the first systematic exploration of aligning LLMs with judgments. 2) We introduce a novel framework, CUT, that facilitates the alignment of LLMs through fine-grained inappropriate content detection and correction based on judgments. 3) Our results showcase the effectiveness of CUT in aligning LLMs across cold-start and warm-start scenarios, generalist and specialist applications, as well as offline and online settings. 4) Our analysis indicates that judgments hold greater potential over rewards for aligning LLMs.
@@ 2 Related Work
Prompting. Prompting does not alter the parameters of LLMs. Instead, it leverages judgments on previous responses to elicit improved responses from LLMs (Welleck et al., 2022; Akyurek et al., 2023). Judgments can be sourced from diverse aspects (Nathani et al., 2023) and the refinement process can be iterated multiple times (Yang et al., 2022; Peng et al., 2023a; Madaan et al., 2023). However, these methods consume more computation than single-pass generation and usually rely on the in-context learning capabilities of the LLMs (Brown et al., 2020; Liu et al., 2023b).
Fine-tuning. Fine-tuning aims to train a better LLM directly. Scalar rewards have been extensively used through RL, particularly PPO (Schulman et al., 2017; Ziegler et al., 2019; Yang et al., 2023). However, these RL approaches are notoriously unstable and complex (Zheng et al., 2023). To stabilize RL, Ramamurthy et al. (2023) propose to reduce the action space through truncation and Peng et al. (2023b) employ an advantage model and selective rehearsal. In addition, many studies aim to design simpler alternatives to RL. Dong et al. (2023); Touvron et al. (2023) treat rewards as a ranking criterion and simply train models using the best model-generated responses. There are also attempts to leverage the results of prompting for training a better model. That is, the improved response elicited by the judgment is employed as new training data (Scheurer et al., 2022, 2023; Yu et al., 2023). However, these methods still suffer from the incapability to learn from mistakes. Rafailov et al. (2023); Yuan et al. (2023); Song et al. (2023); Hong et al. (2024) demonstrate that LLMs themselves can be used as reward models and derive different training objectives to eliminate the need for RL. Zhang et al. (2023); Liu et al. (2023a) relabel the input using the reward received by the response, referred to as Hindsight. Hindsight allows LLMs to generate responses of different qualities.
@@ 3 Preliminaries
In this section, we first lay out a formal problem definition of aligning LLMs with judgments and then present a survey of three potential methods that can be adapted for tackling this problem.
@@ 3.1 Problem Setting
Suppose that there is a set of instruction-response-judgment triplets $(\bm{x},\bm{y},\bm{j})$, where the instruction $\bm{x}=[x_{1},\ldots,x_{M}]$, the response $\bm{y}=[y_{1},\ldots,y_{N}]$, and the judgment $\bm{j}=[j_{1},\ldots,j_{Q}]$ are token sequences of length $M$, $N$, and $Q$, respectively. The response may exhibit flaws or be considered entirely satisfactory. The judgment provides an analysis of the strengths and weaknesses of the response, which can be drafted either by humans or AI models (Akyurek et al., 2023; Li et al., 2023). The goal of aligning LLMs with judgments is to enable LLMs to retain appropriate behaviors mentioned in the strengths, and more importantly, address the weaknesses to prevent future misbehavior.
Depending on whether the responses $\bm{y}$ are from the LLM to be aligned, the learning process can be classified into two distinct types: offline alignment and online alignment. In offline alignment, the target LLM learns from an off-the-shelf, model-agnostic dataset. Conversely, in online alignment, the target LLM reflects on its own outputs through direct interactions with a judge. This online alignment process can be conducted iteratively, akin to how humans continuously improve their skills by receiving ongoing feedback from others over time.
Instruction: $\bm{x}$ Response: $\bm{y}$ Judgment: $\bm{j}$ $\bm{x}\xrightarrow{}\bm{y}$ $[\bm{x},\bm{j}]\xrightarrow{}\bm{y}$
@@
Each pack was 5 pounds and it cost 5.50. So 5 * 5.50 = \$27.50.
@@
Each pack was 5 pounds and it cost 5.50. So 5 * 5.50 = \$27.50.
@@
Table 1: The illustration of three categories of alignment data. $\bm{x}\xrightarrow{}\bm{y}$ and $[\bm{x},\bm{j}]\xrightarrow{}\bm{y}$ indicate if the response aligns with the instruction or the combination of instruction and judgment, respectively.
3.2 Potential Solutions
Forward Prediction refers to sequentially predicting the response and its judgment (Chen et al., 2024), which was originally proposed in dialogue generation (Weston, 2016; Li et al., 2017). It can be seamlessly adapted to our problem. Specifically, the LLM is trained with the maximum likelihood estimation (MLE) objective to first generate the response $\bm{y}$ based on the instruction $\bm{x}$ and subsequently generate the judgment $\bm{j}$ based on the combined sequence $[\bm{x},\bm{y}]$.
$L_{f}=-\frac{1}{N}\sum_{t}\log p(y_{t}|y_{<t},\bm{x})-\frac{1}{Q}\sum_{t}\log p(j_{t}|j_{<t},\bm{y},\bm{x})$ (1)
Imitation learning from language feedback (ILF) asks the LLM to refine the initial response $\bm{y}$ based on the feedback $\bm{j}$ to be an improved response $\hat{\bm{y}}$.
$\hat{\bm{y}}=\textbf{LLM}(\bm{x},\bm{y},\bm{j})$ (2)
The improved response $\hat{\bm{y}}$ can be directly paired with the initial instruction $\bm{x}$ to fine-tune the LLM under the MLE objective (Bai et al., 2022b; Scheurer et al., 2022, 2023).
$L_{i}=-\frac{1}{N}\sum_{t}\log p(\hat{y}_{t}|\hat{y}_{<t},\bm{x})$ (3)
Hindsight rewrites the instruction $\bm{x}$ based on the scalar rewards received by the response $\bm{y}$ (Zhang et al., 2023; Liu et al., 2023a). For instance, if a response receives a scalar reward below a certain threshold, the phrase “generate a good answer" is appended to the original instruction. This approach can be naturally extended to our problem setting. Concretely, the LLM is trained to generate the response $\bm{y}$ conditioned on the sequence $[\bm{x},\bm{j}]$.
$L_{h}=-\frac{1}{N}\sum_{t}\log p(y_{t}|y_{<t},\bm{x},\bm{j})$ (4)
However, in Forward Prediction, learning to generate judgments does not necessarily translate into enhanced response generation, given that response generation precedes judgment generation. The indirect usage of judgment in ILF limits its capacity to spot and rectify weaknesses underscored in judgments. Hindsight employs unsatisfactory responses as MLE targets, which inevitably increases the risk of generating unsatisfactory responses. In summary, we contend that existing methods cannot fully capitalize on judgments, which motivates us to design a better solution.
4 Contrastive Unlikelihood Training
To overcome the limitations mentioned in § 3, we propose CUT, a fine-tuning framework to align LLMs with judgments. The core idea of CUT is summarized as Learning from Contrasting. We contrast the response generation under different conditions to shed light on the appropriate behavior that the LLM should keep, as well as the specific content necessitating adjustments. Based on these insights, we use MLE for appropriate content and UT (Welleck et al., 2020) for inappropriate content.